### v3

# Non-minimally Coupled Running Curvaton for DESI-preferred Dynamical Dark Energy and Hubble Tension

Abstract
Recent DESI 2025 results have renewed interest in dynamical dark energy models with phantom crossing. In this work, we extend the running curvaton scenario by introducing a non-minimal gravitational coupling, $\xi\chi^{2}R$. We show that this coupling can drive the late-time equation of state into the DESI-preferred region in the $(w_{0},w_{a})$ plane. We further find that, at leading order, the early-universe predictions of the original running curvaton model can be preserved through a parameter redefinition, while the scalar sector remains free of gradient instability in the regime considered. A parameter scan identifies viable regions and a benchmark solution that also shifts the background-inferred value of $H_{0}$ upward. These results suggest that the non-minimally coupled running curvaton provides a viable framework for DESI-motivated dark energy evolution, with possible implications for the Hubble tension.
I Introduction
[…]
Specifically, the SH0ES collaboration reported a local value of $H_{0}=73.04\pm 1.04$ km/s/Mpc Riess et al. (2022), leading to a tension exceeding $5\sigma$. This discrepancy is further reinforced by independent late-time probes, including strong gravitational lensing from the H0LiCOW/TDCOSMO collaboration Wong et al. (2020) and the TRGB (Tip of the Red Giant Branch) method Freedman (2021), although some debates on systematic uncertainties remain. This ”Early vs. Late” dichotomy strongly suggests the need for physics beyond the standard paradigm, potentially involving either early-time modifications to the sound horizon Poulin et al. (2019) or late-time dynamical dark energy evolution. Most notably, recent results from the Dark Energy Spectroscopic Instrument (DESI) 2025, when combined with Supernovae Type Ia (SNIa) and CMB datasets, favor a dynamical dark energy component over a cosmological constant Gu et al. (2025); Wang (2024). Specifically, the data prefers an equation of state that crosses the so-called “phantom divide” (w<1w<-1) at late times, characterized by the parameter space $w_{0}>-1$ and $w_{a}<0$ Gu et al. (2025); Rubin et al. (2023); Brout et al. (2022); Abbott et al. (2024); Colgáin et al. (2024); Pan and Ye (2025).
This observational hint poses a significant theoretical challenge. Canonical single-field quintessence models are typically restricted to the region $w\geq-1$ Caldwell et al. (1998); Zlatev et al. (1999), while naive attempts to realize phantom behavior often lead to pathologies such as ghost degrees of freedom or gradient instabilities Vikman (2005); Hu (2005); Caldwell et al. (2003). Consequently, constructing a consistent theoretical framework that explains both the early inflationary epoch and the specific late-time phantom dynamics favored by DESI—without introducing instabilities—remains an open problem Creminelli et al. (2009); Feng et al. (2005); Nojiri et al. (2005). However, this issue can be addressed by introducing a non-minimal coupling Yao et al. (2025); Wolf et al. (2026); Tiwari et al. (2025); Sánchez López et al. (2025); Wang et al. (2025); Tita et al. (2025); Wolf et al. (2025); Adam et al. (2025). Although numerous approaches exist to study evolving dark energy—including the Quintom scenario Cai et al. (2025); Yang et al. (2025, 2024), $f(R)$ gravitational models Nojiri et al. (2025a, b); Odintsov and Oikonomou (2026), even using the minimally coupling to investigate the dynamical DE Gialamas et al. (2025a) that represents the local effects of DE Gialamas et al. (2025b). We restrict our attention in this paper to the non-minimally coupled dark energy model.
Under the framework of non-minimally coupling, there are some models for investigating the Hubble tension. The $f(R)$ gravity could account for this problem in the Jordan frame Schiavone et al. (2023); Montani et al. (2026, 2023); Fazzari et al. (2025). Additionally, Ref. Barroso Varela and Bertolami (2024) demonstrated that the non-minimally coupling between the curvature and gravity could confront with the CMB, without the traditional cosmological constant. Meanwhile, this puzzle could also be released within $2\sigma$ in the Horndeski gravity Tiwari et al. (2024). The most popular method for explaining the evolving dark energy is the Quintessence Sánchez López et al. (2025), that could also address this problem using the newest DESI data.
The “Running Curvaton” model Liu and Xu (2020a) was previously proposed as a unified approach to connect early-universe physics with late-time acceleration. In this framework, the curvaton field generates primordial perturbations during inflation and survives to serve as the dark energy candidate today. However, in its original minimally coupled form, the model behaves as a thawing quintessence field, which is theoretically unable to access the phantom regime (w<1w<-1) suggested by the new data. In this paper, we propose an extension of the Running Curvaton model by introducing a non-minimal gravitational coupling of the form $\xi\chi^{2}R$. We demonstrate that this modification provides the necessary geometric corrections to the effective energy-momentum tensor in the Jordan frame, enabling the equation of state to cross the phantom divide naturally Faraoni (2000); Elizalde et al. (2004); Gannouji et al. (2006). This extension allows the model to populate the $w_{0}-w_{a}$ observational plane favored by DESI 2025, resolving the tension faced by the standard model.
A critical concern when modifying a unified model is whether the new physics spoils successful predictions for the early universe. We address this by deriving a parameter re-tuning scheme. We show that the geometric correction to the effective mass of the curvaton during inflation can be exactly compensated by shifting the coupling parameters. As a result, the standard predictions for the spectral index ($n_{s}$) and non-Gaussianity ($f_{NL}\approx 5/4r_{dec}$) are strictly preserved, ensuring consistency with Planck observations Akrami et al. (2020).
Furthermore, we rigorously check the stability of the model. By mapping our action to the general Horndeski scalar-tensor theory Horndeski (1974); Deffayet et al. (2011); Kobayashi et al. (2011), we confirm that the sound speed of scalar perturbations remains positive ($c_{s}^{2}=1$), guaranteeing that the model is free from gradient instabilities despite the effective phantom equation of state. The local effect of modified gravity is screened due to the high-density environment of the Solar System Ye et al. (2025).
The paper is organized as follows: Section II outlines the non-minimally coupled Running Curvaton model. Section III demonstrates the preservation of early-universe perturbation predictions via parameter re-tuning. Section IV analyzes the late-time evolution, deriving the effective equation of state and confronting the model with DESI constraints and stability requirements, meanwhile it will automatically account for the Hubble tension. Finally, Section V concludes with a discussion of our findings.
II The Model
We consider the Running Curvaton field $\chi$ non-minimally coupled to gravity. The action in the Jordan frame is given by:
$S=\int d^{4}x\sqrt{-g}\left[\frac{1}{2}\left(M_{P}^{2}-\xi\chi^{2}\right)R-\frac{1}{2}(\nabla\chi)^{2}-V(\chi)\right]+S_{m},$ (1)
where $M_{P}$ denotes the reduced Planck mass, $\xi$ is the coupling constant, and $S_{m}$ describes the matter sector (comprising the inflaton in the early epoch and standard matter at late times). In this work, we adopt the running curvaton framework proposed in Ref. Liu and Xu (2020b). Prior to presenting our detailed calculations, we outline the essential features of this model. Recall that in the running curvaton scenario, the curvaton mass evolves due to an explicit coupling between the inflaton and the curvaton, as shown in the potential
$V(\chi)=\frac{1}{2}\frac{g}{M_{P}^{2}}V(\phi)\chi^{2}+V_{1}(1-V_{0}e^{-\lambda\chi/M_{P}}),$ (2)
where $\phi$ denotes the inflaton, while $V_{1}$, $V_{0}$, and $\lambda$ are parameters constrained by the DESI observations. The first term represents the coupling, which originates from parametric resonance during preheating Kofman et al. (1994, 1997), as discussed in our previous work Liu and Xu (2020b). This term dominates over the second term, meaning the effective mass is primarily determined by the first part. Consequently, the second part is negligible compared to the first, which we treat as fixed in the following analysis. After the inflaton decays, the first term of the potential (2) vanishes. However, the second term survives, approaching a constant value of the same order as the cosmological constant.
Motivation for the modified potential.
The choice of the late-time potential in Eq. (2) is motivated by phenomenology rather than by an attempt to claim a full statistical fit. In the original minimally coupled Running Curvaton setup, the pure exponential potential leads to late-time dynamics similar to thawing quintessenc Wolf and Ferreira (2023), whose trajectories in the $(w_{0},w_{a})$ plane are in tension with, or at best only marginally consistent with, the region preferred by recent DESI data Gu et al. (2025); Abdul Karim et al. (2025). By contrast, the modified form
$V(\chi)=V_{1}\left(1-V_{0}e^{-\lambda\chi/M_{P}}\right)$
introduces additional shape freedom that allows the field evolution to interpolate more efficiently between a nearly frozen stage and a later rolling phase. When combined with the non-minimal coupling $\xi\chi^{2}R$, this structure more readily generates trajectories exhibiting the past phantom crossing behavior favored by the DESI-motivated region with $w_{0}>-1$ and $w_{a}<0$ Wolf et al. (2024). Therefore, the potential in Eq. (2) should be understood as a phenomenologically motivated extension that facilitates access to the desired late-time dynamics in our parameter scan, rather than as a uniquely selected or fully fitted form.
The investigations above provide a microscopic origin for dark energy (DE). Drawing on our previous work Liu and Xu (2020b), we summarize the complete picture for DE as follows. During the preheating epoch, the inflaton decays into fundamental particles, such as Higgs bosons and fermions. Simultaneously, a portion of the inflaton energy is transferred to the curvaton via their coupling; this curvaton subsequently generates the curvature perturbations observed in the CMB. Because the curvaton is a light, long-lived field compared to the inflaton, it survives into the late-time universe where the second term of the potential (2) eventually dominates. Due to the non-minimal coupling, the DE equation of state can cross the phantom divide (w=1w=-1). In the following sections, we demonstrate that this model is consistent with the original running curvaton framework Liu and Xu (2020b).
III EARLY-TIME PERTURBATIONS: PRESERVING PREDICTIONS
In the following calculations, we utilize the Friedmann-Lemaître-Robertson-Walker (FLRW) background metric:
$ds^{2}=-dt^{2}+a^{2}(t)\delta_{ij}dx^{i}dx^{j},$ (3)
where $a(t)$ is the scale factor. A primary motivation for the running curvaton model is its ability to generate the observed curvature perturbations. A key concern, however, is whether the introduction of the non-minimal coupling $\xi\chi^{2}R$ spoils these successful early-universe predictions. In this section, we demonstrate that the model’s consistency is strictly preserved through a straightforward parameter re-tuning. To implement this, we first derive the effective mass of the curvaton.
III.1 Modified Effective Mass
To derive the effective mass of the curvaton field during inflation, we begin with the Jordan frame action in Eq. (1). The effective curvaton mass can be defined as $m_{\rm eff}^{cur}=V^{\prime\prime}(\bar{\chi})$; alternatively, it can be extracted from the equation of motion (EOM) for the curvaton perturbation. We define the curvaton field as $\chi(x_{\mu})=\bar{\chi}(t)+\delta\chi(x_{\mu})$, where $\bar{\chi}$ is the homogeneous background and $\delta\chi$ is the perturbation. By substituting this decomposition into the action (1) and expanding to second order in $\delta\chi$, we obtain the action for the curvaton perturbations:
$S_{\delta\chi}=\int d^{4}x\sqrt{-g}\left[-\frac{1}{2}g^{\mu\nu}\partial_{\mu}\delta\chi\partial_{\nu}\delta\chi-\frac{1}{2}(V^{\prime\prime}+\xi R)\delta\chi^{2}\right].$ (4)
Varying with actionm (4), we could obtain the EOM of curvaton perturbation of curvaton as follows,
$\frac{1}{\sqrt{-g}}\partial_{\mu}(\sqrt{-g}g^{\mu\nu}\partial_{\nu}\delta\chi)-(V^{\prime\prime}(\bar{\chi})+\xi R)\delta\chi=0\,.$ (5)
Making use of the background metric (3), we could obtain
$\delta\ddot{\chi}+3H\delta\dot{\chi}+k^{2}\delta\chi+\left(V^{\prime\prime}(\bar{\chi})+\xi R\right)\delta\chi=0\,$, (6)
where we could define effective mass
$m_{\rm eff}^{2}=m_{\rm orig}^{2}+\xi R\,$ (7)
where there is geometric correction for the effective mass of curvaton. Based on FLRM metric (3), the Ricci scalar can be easily obtained as
$R=6(2H^{2}+\dot{H})\,$, (8)
where $H=\frac{\dot{a}}{a}$. In light of slow-roll condition, we have $\epsilon=-\frac{\dot{H}}{H^{2}}\ll 1$. Therefore, the term of $12H^{2}$ is dominant compared to $6\dot{H}$. The Ricci scalar can be simplified into
$R\approx 12H^{2}.$ (9)
Finally, the effective mass of curvaton is explicitly obtained as follows,
$m_{\rm eff}^{2}=m_{\rm orig}^{2}+12\xi H^{2}\,.$ (10)
From the original Liu and Xu (2020b), the original effective mass in earl universe for curvaton is determined by the coupling part of potential
$m_{\rm orig}\approx\frac{g}{M_{P}^{2}}m^{2}\phi^{2}$ (11)
In this calculation, we assume a simple quadratic inflationary potential, $V(\phi)=\frac{1}{2}m_{\phi}^{2}\phi^{2}$. As indicated by the curvaton potential (2), our framework is versatile enough to incorporate multiple curvaton modes. A distinct advantage of the curvaton mechanism is that it relaxes the constraints on the inflaton potential, removing the necessity for a plateau-like shape.
During the early universe, the energy density is primarily dominated by the inflaton potential. Under the slow-roll approximation, the kinetic term is negligible, and the Hubble parameter is given by:
$3M_{P}^{2}H^{2}\approx\frac{1}{2}m_{\phi}^{2}\phi^{2}\implies m_{\phi}^{2}\phi^{2}\approx 6M_{P}^{2}H^{2}.$ (12)
Utilizing this approximation, the original curvaton mass can be expressed in terms of the Hubble parameter as follows:
$m_{\rm orig}^{2}\approx\frac{g_{0}}{M_{P}^{2}}(6M_{P}^{2}H^{2})=6g_{0}H^{2},$ (13)
Finally, the effective mass of curvaton in this work can be denoted by
$m_{\rm eff}^{2}=6g_{0}H^{2}+12\xi H^{2}=6H^{2}(g_{0}+2\xi),$ (14)
This geometric representation to the mass is the key factor that necessitates the re-tuning of the coupling constant discussed in the following subsection.
III.2 Re-tuning Scheme
Based on the detailed calculation in Appendix A and based on the effective mass of curvaton (10), we could easily derive the spectral index of curvaton
$n_{\chi}-1=4(g_{0}+2\xi)=4g_{0}^{obs},$ (15)
where we have used the de-Sitter approximation which captures the complete information for curvaton. Compared with the original work Liu and Xu (2020b), where $n_{\chi}^{\rm orig}-1=\frac{2}{3}\frac{m_{\rm orig}^{2}}{H^{2}}$, it explicitly indicates that the $g_{0}+2\xi$ will re-tune in order to maintain the same value of $g_{0}^{\rm orig}$ via Liu and Xu (2020b).
III.3 Non-Gaussianity and Tensor Modes
The local non-Gaussianity parameter $f_{\rm NL}$ in curvaton models is determined by the energy fraction at the time of decay, $r_{\rm dec}\approx\rho_{\chi}/\rho_{\rm tot}$. The energy density of $\rho_{\chi}$ is mainly determined by $\frac{1}{2}m_{\rm eff}\chi^{2}$. Through appendix B, the prediction for non-Gaussianity remains:
$f_{\rm NL}\approx\frac{5}{4r_{\rm dec}}\,$, (16)
which is consistent with Planck constraints for sufficiently large $r_{\rm dec}$. In our model with non-minimal coupling $\xi\chi^{2}R$, this derivation holds valid due to two reasons:
  • Re-tuning Mechanism: The parameter re-tuning ($g_{0}+2\xi=g_{0}^{obs}$) ensures that the field fluctuations $\delta\chi$ and background value $\bar{\chi}$ during inflation are identical to the standard minimally coupled case.
  • Radiation Domination: During the decay epoch, the universe is radiation-dominated, implying $R\approx 0$. Thus, the non-minimal coupling term $\xi R\chi^{2}$ vanishes, and the energy density $\rho_{\chi}$ follows the standard canonical form used above.
Consequently, the prediction $f_{NL}\approx\frac{5}{4r_{dec}}$ is preserved.
Finally, regarding tensor perturbations, Appendix C has give the detailed calculation to the tensor mode as shown in Eq. (76). The non-minimal coupling introduces a correction of order $\mathcal{O}(\chi^{2}/M_{P}^{2})$ to the tensor power spectrum. Since the curvaton is sub-dominant during inflation, the standard inflationary predictions for gravitational waves (and the tensor-to-scalar ratio $r$) are preserved to high accuracy.
In this section, we generally discussed the constraint from the universe. Due to the re-tuning parameter $g_{0}^{obs}=g_{0}+2\xi$, the Spectral index, Non-Gaussanity and tensor model are all consistent with CMB observations.
IV Late-Time Evolution and Constraints
In this section, we will mainly focus on the evolution of DE originated from curvaton. We need to investigate the stability for our model (1) since many DE models suffer the gradient instabilities Vikman (2005); Creminelli et al. (2009); Gubitosi et al. (2013); Wolf and Lagos (2019); Creminelli et al. (2020). For us, we expect our model will not suffer this instability problem. Thus, the stability based on the sound speed is of particular importance.
IV.1 Stability Analysis: Sound Speed
A crucial consistency check for any dark energy model crossing the phantom divide is the stability of scalar perturbations. Specifically, we must ensure that the effective sound speed squared, $c_{s}^{2}$, remains positive to avoid gradient instabilities.
Our action (Eq. 1) falls within the general Horndeski scalar-tensor theory Horndeski (1974), whose Lagrangian is defined by the functions $G_{i}(\chi,X)$, where $X\equiv-\frac{1}{2}\nabla_{\mu}\chi\nabla^{\mu}\chi$ is the kinetic term. Mapping our non-minimally coupled model to the Horndeski framework, we identify:
$G_{2}(\chi,X)$ $=X-V(\chi)\,$, (17)
$G_{4}(\chi,X)$ $=\frac{1}{2}(M_{P}^{2}-\xi\chi^{2})\,$, (18)
$G_{3}(\chi,X)$ $=G_{5}(\chi,X)=0\,.$ (19)
Note that our kinetic term is canonical, implying $G_{2,X}=1$ and $G_{2,XX}=0$. Furthermore, the non-minimal coupling function $G_{4}$ depends solely on the field value $\chi$, so $G_{4,X}=0$.
The effective sound speed squared for scalar perturbations in Horndeski theory is given by $c_{s}^{2}=\mathcal{F}_{S}/\mathcal{G}_{S}$, where $\mathcal{F}_{S}$ and $\mathcal{G}_{S}$ are determined by the background evolution. In the absence of derivative couplings (i.e., $G_{4,X}=0$ and $G_{3}=G_{5}=0$), these expressions simplify significantly:
$\mathcal{F}_{S}$ $=2G_{2,X}+\dots=2\,$, (20)
$\mathcal{G}_{S}$ $=2G_{2,X}+4XG_{2,XX}+\dots=2\,.$ (21)
Consequently, the sound speed is:
$c_{s}^{2}=\frac{\mathcal{F}_{S}}{\mathcal{G}_{S}}=1\,.$ (22)
This result ($c_{s}^{2}=1>0$) confirms that despite the effective equation of state crossing into the phantom regime ($w_{\rm eff}<-1$) due to the non-minimal coupling $\xi\chi^{2}R$, the scalar field perturbations propagate at the speed of light and are free from gradient instabilities. The phantom behavior in our model is a frame-dependent effect arising from the modified Einstein equations, rather than a pathological kinetic term. In addition to the positivity of $c_{s}^{2}$, viability of the model also requires the effective Planck mass squared, $M_{\rm eff}=M_{P}^{2}-\xi R^{2}$, to remain positive throughout the cosmological evolution considered here. Futhermore, the tensor propagation speed remains luminal $c_{T}^{2}=1$ due to $G_{5}=0$ and $G_{4,X}=0$, in agreement with current gravitational-wave constraints.
IV.2 Remarks on Local Screening and Possible Solar-System Viability
Due to the environmental dependence of the scalar field’s dynamics, a feature inherent to non-minimally coupled models, the effective mass squared of the curvaton field is given by Eq. (7):
$m_{\text{eff}}^{2}(\chi)=V^{\prime\prime}(\chi)+\xi R.$ (23)
In the high-density environment of the Solar System, the trace of the Einstein equations dictates that the Ricci scalar tracks the local matter density, $R\approx\rho_{\text{loc}}/M_{P}^{2}$. This induces a geometric correction to the mass:
$m_{\text{loc}}^{2}\approx V^{\prime\prime}(\chi)+\xi\frac{\rho_{\text{loc}}}{M_{P}^{2}}.$ (24)
For coupling strengths of order unity ($\xi\sim\mathcal{O}(1)$), the high density of local matter ($\rho_{\text{loc}}\gg\rho_{\text{cosmo}}$) generates a large effective mass for the scalar field locally. This phenomenology is analogous to the Symmetron screening mechanism Hinterbichler and Khoury (2010); Khoury (2010), where the field is pinned to a near-zero value in high-density regions, or the Chameleon mechanism Khoury and Weltman (2004), where the field becomes sufficiently heavy to suppress long-range interactions.
Consequently, the interaction range of the fifth force, $\lambda\sim m_{\text{loc}}^{-1}$, becomes significantly smaller than the scale of Solar System experiments ($\lambda\ll\text{AU}$). The scalar interaction is Yukawa-suppressed by a factor of $e^{-m_{\text{loc}}r}$, effectively restoring General Relativity ($\gamma\approx 1$) locally. This screening mechanism allows the model to satisfy local gravity constraints without requiring the cosmological field value $\chi_{0}$, which drives the late-time phantom crossing favored by DESI Ye et al. (2024), to be negligible.
Here, we emphasize that this argument suggests a possible suppression of fifth-force effects in high-density environments. A quantitative confrontation with Solar-System tests, however, would require a dedicated analysis of the static spherically symmetric solution and the corresponding post-Newtonian parameters, which lies beyond the scope of the present work.
IV.3 Modified Friedmann Equations
In a flat FLRW universe, the variation of Action (1) yields the modified Friedmann equations (Detailed calculation can be found in Appendix D). Neglecting the inflaton field at late times, we have:
$3(M_{P}^{2}-\xi\chi^{2})H^{2}=\rho_{m}+\frac{1}{2}\dot{\chi}^{2}+V(\chi)+6\xi H\chi\dot{\chi},$ (25)
and
$-2(M_{P}^{2}-\xi\chi^{2})\dot{H}$ $=$ $\rho_{m}+\dot{\chi}^{2}+2\xi H\chi\dot{\chi}$ (26)
- $2\xi(\dot{\chi}^{2}+\chi\ddot{\chi}).$
IV.4 Effective Equation of State
To compare with observations, we cast the equations into the standard Einstein form: $3M_{P}^{2}H^{2}=\rho_{m}+\rho_{DE}^{eff}$. The effective dark energy density and pressure in the Jordan frame are defined as:
$\rho_{DE}^{eff}=\frac{1}{2}\dot{\chi}^{2}+V(\chi)+6\xi H\chi\dot{\chi}+3\xi\chi^{2}H^{2},$ (27)
$p_{DE}^{eff}$ $=$ $\frac{1}{2}\dot{\chi}^{2}-V(\chi)-2\xi(\chi\ddot{\chi}+\dot{\chi}^{2})$ (28)
- $4\xi H\chi\dot{\chi}-\xi\chi^{2}(2\dot{H}+3H^{2}).$
The effective equation of state is defined as $w_{\rm eff}=p_{DE}^{\rm eff}/\rho_{DE}^{\rm eff}$. To investigate its evolution numerically, we rewrite the dynamics in terms of redshift $z$. Using the relation $1+z=1/a$, we transform the time derivative to $\dot{\chi}=-(1+z)H\chi^{\prime}$, where the prime denotes differentiation with respect to $z$. Consequently, we derive the Hubble parameter as a function of $z$ as follows:
$H^{2}(z)=\frac{\rho_{m,0}(1+z)^{3}+V(\chi)}{3(M_{P}^{2}-\xi\chi^{2})-\frac{1}{2}(1+z)^{2}\chi^{\prime 2}+6\xi(1+z)\chi\chi^{\prime}},$ (29)
where $\rho_{m,0}$ is the present value of matter’s density. Similarly, we also could derive the derivative of $H(z)$ as follows,
$H^{\prime}(z)=-\frac{\dot{H}}{(1+z)H}=\frac{S_{1}(z)-2\xi\chi S_{2}(z)}{(1+z)H\Delta},$ (30)
where we have defined $S_{1}(z)=\rho_{m0}(1+z)^{3}+(1-2\xi)(1+z)^{2}H^{2}\chi^{\prime 2}-2\xi(1+z)H^{2}\chi\chi^{\prime}$, $S_{2}(z)=3(1+z)H^{2}\chi^{\prime}-V^{\prime}(\chi)-12\xi H^{2}\chi$ and $\Delta=2(M_{P}^{2}-\xi\chi^{2}+6\xi^{2}\chi^{2})$. Based on eq. (29) and eq. (30), we could numerically simulate the crossing the phantom divide.
IV.5 Crossing the Phantom Divide
While the standard canonical model ($\xi=0$) is constrained to the quintessence regime ($w\geq-1$), the introduction of non-minimal coupling ($\xi\neq 0$) generates geometric terms proportional to $H^{2}$ and $\dot{H}$ that contribute negatively to the effective pressure. Mathematically, achieving a phantom equation of state ($w_{\rm eff}<-1$) necessitates $p_{DE}^{\rm eff}+\rho_{DE}^{\rm eff}<0$. For our specific action, this condition takes the form:
$\rho_{DE}^{eff}+p_{DE}^{eff}=\dot{\chi}^{2}(1-2\xi)+2\xi(H\chi\dot{\chi}-\chi\ddot{\chi}-\chi^{2}\dot{H}).$ (31)
By tuning $\xi$ (e.g., small positive values) and given the dynamics of the rolling scalar field, this quantity can become negative, allowing the model to fit the $w(z)$ behavior suggested by DESI 2025 (crossing from w<1w<-1 in the past to w>1w>-1 today).
To numerically solve the background evolution of the curvaton field, we impose the initial conditions $\dot{\chi}_{ini}=0$ and set $\chi_{ini}$ to a negligible non-zero value. Following Ref. Wolf et al. (2024), we tune the parameters $\{\xi,\lambda,V_{1},V_{0}\}$ to satisfy the equations of motion, ensuring that the resulting equation of state crosses the phantom divide. Regarding units, we adopt the convention of Ref. Ye et al. (2025), where setting $V_{1}\approx 3H_{0}^{2}\Omega_{\varphi}$ provides a robust basis for simulating the dark energy evolution with respect to redshift $z$. Throughout this work, we set $M_{P}=1$, and express $V_{0}$ in units of $M_{P}^{2}H_{0}/h$. The parameters $\lambda$ and $\chi$ are dimensionless.
Figure 1: Evolution of the effective dark-energy equation of state $w(z)$ for four representative parameter choices in the non-minimally coupled running curvaton model. The curves illustrate that increasing $\xi$ facilitates access to a phantom phase at intermediate redshift, while the trajectories approach w=1w=-1 as $\xi$ approaches zero.
Based on Eqs. (29) and (30), we numerically simulate the redshift dependence of the effective equation of state. As shown in Fig. 1, the non-minimal coupling $\xi$ is the key driver for crossing the phantom divide; specifically, the trajectory approaches the cosmological constant limit ($w\to-1$) as $\xi\to 0$. Analyzing Eqs. (27) and (28) reveals that the term $\xi\chi^{2}(2\dot{H}+3H^{2})$ contributes a negative pressure correction, suppressing $w(z)$ during late times before it rises again due to field evolution. This allows the non-minimally coupled Running Curvaton to resolve the phantom divide problem without pathological kinetic terms. Furthermore, the parameter freedom in $(\lambda,\chi,V_{0},V_{1})$ allows the model to be mapped extensively onto the $w_{0}-w_{a}$ observational plane.
IV.6 Comparison with DESI-preferred Regions
Our goal in this section is not to perform a full global fit, but rather to identify regions of parameter space in which the model can reproduce the DESI-preferred late-time dark-energy behavior.
In this section, we will show our model will be consistent with DESI Gu et al. (2025). In our case, $w_{0}$ denotes the present value of $w_{\rm eff}|_{a=1}$. Based on (27) and (28), we could derive
$w_{0}=w_{\text{eff}}(a)|_{a=1}$
$=\left.\frac{\frac{1}{2}\dot{\chi}^{2}-V(\chi)-2\xi(\chi\ddot{\chi}+\dot{\chi}^{2})-4\xi H\chi\dot{\chi}-\xi\chi^{2}(2\dot{H}+3H^{2})}{\frac{1}{2}\dot{\chi}^{2}+V(\chi)+6\xi H\chi\dot{\chi}+3\xi\chi^{2}H^{2}}\right|_{z=0},$ (32)
where $w_{\text{eff}}(a)=\frac{p_{\text{DE}}^{\text{eff}}}{\rho_{\text{DE}}^{\text{eff}}}$. As for the formula of $w(a)$, it is defined in terms of redshift as follows,
$w_{a}=\left.\frac{dw_{\text{eff}}}{dz}\right|_{z=0}.$ (33)
Being armed with these two formulas, we could numerically simulate the $w$.
[…]
Figure 2: Representative model trajectories in the $w_{0}-w_{a}$ plane compared with the DESI-motivated confidence contours shown for reference. The colored markers correspond to the four benchmark parameter choices used in Fig. 1. This figure is intended as a visual comparison.
[…]
Figure 3: Distributions of accepted parameter samples for the non-minimally coupled running curvaton model, obtained from the parameter scan described in Sec. IV F. The darker and lighter shaded regions indicate the regions enclosing $68\%$ and $95\%$ of the accepted samples, respectively.
To constrain the free parameters of the non-minimal coupling model, namely $\theta=\{\xi,\lambda,V_{0},V_{1}\}$, we performed a Monte Carlo parameter scan using the rejection sampling technique. We assumed uniform priors for all parameters within physically motivated ranges: $\xi\in[0,5]$, $\lambda\in[0,2]$, $V_{0}\in[0.5,1.5]$, and $V_{1}\in[1.0,6.0]$. For each randomly sampled set of parameters, we numerically solved the background equations of motion (Eq. 32 and Eq. 33) to derive the corresponding dark energy equation-of-state parameters at present day, $w_{0}$, and its evolution, $w_{a}$. These theoretical predictions were then mapped onto the observational $w_{0}-w_{a}$ plane. We utilized the $95\%$ confidence level (C.L.) contours obtained from the combination of DESI DR2 BAO, Pantheon+ supernovae, and Planck CMB data as our acceptance criterion. Specifically, we adopted a top-hat likelihood approach: a parameter set was accepted only if its predicted $(w_{0},w_{a})$ fell within the observational $95\%$ C.L. region; otherwise, it was rejected. Fig. 3 shows the resulting distributions of accepted samples and 2D confidence regions for the model parameters, derived from $20,000$ Monte Carlo trials. Based on the results shown in this figure, the best-fit parameters are $(\xi,\lambda,V_{0},V_{1})=(2.9,0.96,0.76,3.5)$. We observe that Model B lies almost entirely outside the DESI constraints, as its $\xi$ value is smaller than the minimum threshold of $\xi=1.6$ shown in Fig. 2. In contrast, Model C aligns well with the best-fit region in Fig. 3. Overall, the results in Fig. 3 demonstrate strong consistency with those presented in Fig. 2.
IV.7 Implications for the Hubble Tension
Besides fitting the DESI-preferred dynamical dark energy region in the $w_{0}$$w_{a}$ plane, the late-time phantom crossing in the non-minimally coupled Running Curvaton model also has a direct geometric implication for the Hubble tension. The key point is that the non-minimal coupling modifies the late-time expansion history while leaving the pre-recombination physics essentially unchanged through the re-tuning mechanism
$g_{0}^{\rm obs}=g_{0}+2\xi.$ (34)
As discussed in Sec. III, this guarantees that the primordial perturbation sector is preserved, so that the standard early-universe quantities, including the comoving sound horizon at recombination, remain unchanged to a very good approximation.
IV.7.1 CMB acoustic-scale consistency
The strongest early-universe constraint comes from the precisely measured CMB acoustic scale,
$\theta_{*}=\frac{r_{s}(z_{*})}{D_{M}(z_{*})},$ (35)
where $r_{s}(z_{*})$ is the comoving sound horizon at recombination and $D_{M}(z_{*})$ is the comoving angular-diameter distance to the last-scattering surface. Because the re-tuning mechanism preserves the pre-recombination expansion history in our setup, $r_{s}(z_{*})$ is fixed at essentially the same value as in the standard scenario. Therefore, consistency with the observed $\theta_{*}$ requires that
$D_{M}(z_{*})=c\int_{0}^{z_{*}}\frac{dz}{H(z)}$ (36)
remains unchanged.
It is convenient to rewrite the Hubble parameter in the form
$H(z)=H_{0}E(z),$ (37)
with
$E^{2}(z)=\Omega_{m,0}(1+z)^{3}+\Omega_{{\rm DE},0}^{\rm eff}\,f(z),$ (38)
where the dark-energy evolution factor is
$f(z)=\exp\left[3\int_{0}^{z}\frac{1+w_{\rm eff}(z^{\prime})}{1+z^{\prime}}\,dz^{\prime}\right].$ (39)
Hence,
$D_{M}(z_{*})=\frac{c}{H_{0}}\int_{0}^{z_{*}}\frac{dz}{E(z)}.$ (40)
IV.7.2 Late-time phantom crossing and the shift of $H_{0}$
[…]
$w_{\rm eff}(z)<-1.$ (41)
In this regime,
$1+w_{\rm eff}(z)<0,$ (42)
and therefore the exponent in $f(z)$ becomes negative, implying
$f_{\rm phantom}(z)<1$ (43)
for the relevant redshift interval. As a consequence, the normalized expansion history satisfies
$E_{\rm phantom}(z)<E_{\Lambda{\rm CDM}}(z)$ (44)
at intermediate redshifts where the phantom behavior is active.
[…]
$\int_{0}^{z_{*}}\frac{dz}{E(z)}$ (45)
increases. To keep the CMB distance $D_{M}(z_{*})$ fixed and preserve the observed acoustic scale $\theta_{*}$, the prefactor $H_{0}^{-1}$ must decrease. Equivalently, the inferred present-day Hubble constant must increase:
$H_{0}^{\rm phantom}>H_{0}^{\Lambda{\rm CDM}}.$ (46)
Therefore, the same late-time phantom-crossing dynamics that improves the fit to the DESI data naturally shifts the preferred value of $H_{0}$ upward, partially alleviating the tension between the CMB-inferred and local distance-ladder determinations.
IV.7.3 Numerical implication for the Hubble constant
To quantify the geometric effect discussed above, we numerically integrate the background evolution from $z=0$ to $z\approx z_{*}\simeq 1090$, while imposing the condition that the comoving angular diameter distance to the last-scattering surface remains fixed. Since the re-tuning mechanism preserves the early-universe physics and keeps the sound horizon $r_{s}(z_{*})$ unchanged, the shift in the inferred Hubble constant is entirely induced by the modified late-time expansion history.
For the benchmark solution near the DESI best-fit region, namely
$(\xi,\lambda,V_{0},V_{1})=(2.9,\,0.96,\,0.76,\,3.5),$ (47)
which corresponds to Model C in Fig. 2 and Fig. 3, the numerical integration shows that the preferred Hubble constant is shifted upward to
$H_{0}\approx 73.94~{\rm km\,s^{-1}\,Mpc^{-1}}.$ (48)
[…]
Fig. 4 illustrates this result explicitly. The black dashed curve and red solid curve represent the Gaussian distributions corresponding to the Planck 2018 and SH0ES 2022 determinations, respectively, while the green dash-dotted curve denotes the theoretical prediction of our non-minimally coupled Running Curvaton model for the benchmark parameter set. One can clearly see that the geometric phantom-crossing effect shifts the model prediction from the Planck value toward the SH0ES region, thus providing a viable late-time mechanism for reconciling the two measurements.
[…]
IV.7.4 Remarks on interpretation
We stress that, at the present stage, the resolution of the Hubble tension discussed here should be understood as a background-level geometric mechanism: the late-time phantom evolution lowers $E(z)$ over part of cosmic history, which in turn pushes the CMB-compatible value of $H_{0}$ upward. A complete statistical assessment requires a joint likelihood analysis using CMB, BAO, and supernova data, together with the exact background solution for the non-minimally coupled curvaton field. Nevertheless, Nevertheless, the qualitative mechanism is clear at the background level and follows directly from the modified late-time expansion history induced by the coupling $\xi\chi^{2}R$.
V Discussion and Conclusion
[…]
(a). Resolution of the Phantom Divide Crossing
The most significant achievement of this model is its ability to naturally explain the dynamical dark energy evolution favored by DESI 2025 combined with CMB and SNIa data. We showed that the interplay between the non-minimal coupling $\xi\chi^{2}R$ and the plateau-like potential $V(\chi)$ generates necessary geometric corrections to the effective energy-momentum tensor. This mechanism drives the effective equation of state $w_{eff}$ to cross the phantom divide (w<1w<-1) in the past and evolve towards the quintessence regime (w>1w>-1) today, precisely populating the observational contour of $w_{0}>-1$ and $w_{a}<0$. In light of the most recent DESI 2025 data Gu et al. (2025), Fig. 2 demonstrates that our model falls well within the observational confidence contours. Additionally, the corresponding best-fit model parameters are illustrated in Fig. 3.
(b). Implication to the Hubble tension
The late-time phantom crossing driven by $\xi\chi^{2}R$ modifies the expansion history by lowering the normalized expansion rate $E(z)$ as shown in Sec. IV.7. To maintain the precisely measured CMB acoustic scale $\theta_{*}$, the inferred Hubble constant is naturally shifted upward to $H_{0}\approx 73.94\text{ km s}^{-1}\text{Mpc}^{-1}$. This result falls within the $1\sigma$ region of the SH0ES 2022 local distance-ladder measurement, indicating that the model has the potential to ease the tension at the level of background geometry.
(b). Preservation of Inflationary Predictions
A critical concern in modified gravity models is the potential disruption of early-universe phenomenology. We addressed this by deriving a robust parameter re-tuning scheme, $g_{0}^{obs}=g_{0}+2\xi$. We demonstrated that the geometric correction to the curvaton’s effective mass during inflation can be exactly compensated by shifting the coupling parameters. Consequently, the standard predictions for the primordial power spectrum (spectral index $n_{s}$) and local-type non-Gaussianity ($f_{NL}\approx 5/4r_{dec}$) remain strictly preserved, ensuring the model’s consistency with Planck precision data.
(c). Theoretical Stability and Viability
By mapping the action to the general Horndeski scalar-tensor theory, we performed a comprehensive stability analysis. We confirmed that the model is free from pathological ghost and gradient instabilities, maintaining a positive sound speed squared ($c_{s}^{2}=1$) despite the effective phantom equation of state. Furthermore, our analysis indicate that the couplings of order $\xi\sim\mathcal{O}(1)$ will not impact the global gravity due to screened mechanism.
In summary, the Non-minimally Coupled Running Curvaton offers a theoretically sound and observationally consistent pathway to unify the physics of the early and late universe. It not only alleviates the current tensions in the dark energy sector but also provides a concrete mechanism for dynamical phantom crossing and Hubble tension without violating fundamental stability conditions. Future high-precision surveys, such as Euclid and LSST, will further constrain the derivative of the equation of state, offering a decisive test for the specific $w_{0}-w_{a}$ trajectory predicted by this framework. Furthermore, this work could be extended into multi-field frameworks Liu and Prokopec (2021); Zhang et al. (2023) and $f(R)$ gravity theories Liu (2018); Liu et al. (2018, 2021). The running curvaton model also provides a mechanism for generating primordial black holes (PBHs) as candidates for dark matter Liu (2023). Additionally, one could investigate the effects of non-minimal coupling on PBH formation.
VI ACKNOWLEDGMENTS
LH, BC are supported by the National Natural Science Foundation of China (Grant No. 12165009) and the Hunan Natural Science Foundation (Grants No. 2023JJ30487 and No. 2022JJ40340).
Appendix A Derivation of the Spectral Index with Non-minimal Coupling
In this appendix, we provide a detailed derivation of the spectral index $n_{\chi}$ for the curvaton field non-minimally coupled to gravity. We show that the geometric correction to the effective mass leads directly to the modified spectral tilt used in our re-tuning mechanism.
A.1 Perturbation Equation of Motion
Starting from the action in the Jordan frame as given in (1), we consider the linear perturbations of the curvaton field $\chi(t,\mathbf{x})=\bar{\chi}(t)+\delta\chi(t,\mathbf{x})$. In the flat FLRW background $ds^{2}=-dt^{2}+a^{2}(t)d\mathbf{x}^{2}$, the equation of motion for the Fourier mode $\delta\chi_{k}$ is given by:
$\ddot{\delta\chi}_{k}+3H\dot{\delta\chi}_{k}+\left(\frac{k^{2}}{a^{2}}+m_{\rm eff}^{2}\right)\delta\chi_{k}=0,$ (49)
where the effective mass squared $m_{\rm eff}^{2}$ includes the geometric contribution from the non-minimal coupling:
$m_{\rm eff}^{2}\equiv V^{\prime\prime}(\bar{\chi})+\xi R.$ (50)
During slow-roll inflation, the Ricci scalar is dominated by the Hubble expansion, $R=6(2H^{2}+\dot{H})\approx 12H^{2}$. Thus, the effective mass becomes:
$m_{\rm eff}^{2}\approx V^{\prime\prime}(\bar{\chi})+12\xi H^{2}.$ (51)
A.2 Reduction to Bessel Equation
To solve Eq. (49), we switch to conformal time $\tau=\int dt/a(t)$ and define the canonical variable $u_{k}(\tau)=a(\tau)\delta\chi_{k}(\tau)$. The evolution equation transforms into:
$u_{k}^{\prime\prime}+\left(k^{2}-\frac{a^{\prime\prime}}{a}+a^{2}m_{\rm eff}^{2}\right)u_{k}=0.$ (52)
In the de Sitter limit ($a\approx-1/H\tau$), the effective potential term becomes $(a^{\prime\prime}/a-a^{2}m_{\rm eff}^{2})\approx\frac{1}{\tau^{2}}(2-m_{\rm eff}^{2}/H^{2})$. Consequently, the mode equation takes the form of a Bessel equation:
$u_{k}^{\prime\prime}+\left(k^{2}-\frac{\nu^{2}-1/4}{\tau^{2}}\right)u_{k}=0,$ (53)
where the order $\nu$ is defined by:
$\nu=\sqrt{\frac{9}{4}-\frac{m_{\rm eff}^{2}}{H^{2}}}.$ (54)
A.3 Power Spectrum and Spectral Index
The solution matching the Bunch-Davies vacuum at early times ($|k\tau|\gg 1$) is given by the Hankel function of the first kind:
$u_{k}(\tau)=\frac{\sqrt{\pi}}{2}\sqrt{-\tau}H_{\nu}^{(1)}(-k\tau).$ (55)
On super-horizon scales ($|k\tau|\ll 1$), the asymptotic behavior of the Hankel function yields the power spectrum $\mathcal{P}_{\delta\chi}\equiv\frac{k^{3}}{2\pi^{2}}|\delta\chi_{k}|^{2}$:
$\mathcal{P}_{\delta\chi}(k)\propto k^{3-2\nu}.$ (56)
The spectral index is defined as $n_{\chi}-1\equiv d\ln\mathcal{P}_{\delta\chi}/d\ln k=3-2\nu$. For a light field ($m_{\rm eff}^{2}\ll H^{2}$), we expand $\nu$ to first order:
$\nu\approx\frac{3}{2}\left(1-\frac{2}{3}\frac{m_{\rm eff}^{2}}{3H^{2}}\right)=\frac{3}{2}-\frac{m_{\rm eff}^{2}}{3H^{2}}.$ (57)
Substituting this back into the expression for $n_{\chi}$, we arrive at the standard result modified by the effective mass:
$n_{\chi}-1\approx\frac{2}{3}\frac{m_{\rm eff}^{2}}{H^{2}}.$ (58)
Substituting the explicit form of the mass $m_{\rm eff}^{2}=m_{\rm orig}^{2}+12\xi H^{2}$, we obtain:
$n_{\chi}-1\approx\frac{2}{3}\frac{m_{\rm orig}^{2}}{H^{2}}+8\xi.$ (59)
This linear dependence on $\xi$ necessitates the parameter re-tuning strategy ($g_{0}\to g_{0}-2\xi$) discussed in Section III to preserve consistency with Planck observations.
Appendix B Detailed Calculation of Non-Gaussianity $f_{NL}$
In this appendix, we will derive the specific form of the non-linearity parameter $f_{NL}$.
B.1 Curvature Perturbation Expansion
The curvature perturbation $\zeta$ on uniform density hypersurfaces is related to the energy density fluctuation $\delta\rho$ by:
$\zeta=-H\frac{\delta\rho}{\dot{\rho}}=\frac{1}{3}\frac{\delta\rho}{\rho},$ (60)
assuming the universe is dominated by a fluid with equation of state $w\approx 0$ (oscillating curvaton) or considering the decay into radiation.
For the curvaton field $\chi$, the energy density is given by $\rho_{\chi}\approx m_{\rm eff}^{2}\chi^{2}/2$. The density perturbation, expanded to second order in the field fluctuation $\delta\chi$, is:
$\rho_{\chi}(\bar{\chi}+\delta\chi)$ $=\frac{1}{2}m_{\rm eff}^{2}(\bar{\chi}+\delta\chi)^{2}$
$=\frac{1}{2}m_{\rm eff}^{2}\bar{\chi}^{2}\left(1+2\frac{\delta\chi}{\bar{\chi}}+\left(\frac{\delta\chi}{\bar{\chi}}\right)^{2}\right).$ (61)
Thus, the density contrast is:
$\frac{\delta\rho_{\chi}}{\rho_{\chi}}=2\frac{\delta\chi}{\bar{\chi}}+\left(\frac{\delta\chi}{\bar{\chi}}\right)^{2}.$ (62)
B.2 Transfer to Adiabatic Perturbation
Using the sudden decay approximation, the total curvature perturbation $\zeta$ is a weighted sum of the inflaton and curvaton perturbations. Assuming the inflaton contribution is negligible, $\zeta$ is given by:
$\zeta\approx r_{dec}\zeta_{\chi}=\frac{r_{dec}}{3}\frac{\delta\rho_{\chi}}{\rho_{\chi}},$ (63)
where $r_{dec}\equiv\left.\frac{3\rho_{\chi}}{3\rho_{\chi}+4\rho_{\gamma}}\right|_{decay}\approx\frac{\rho_{\chi}}{\rho_{tot}}$ is the energy fraction of the curvaton at the time of decay (in the limit $r_{dec}\ll 1$).
Substituting the expansion of $\delta\rho_{\chi}/\rho_{\chi}$:
$\zeta=\frac{r_{dec}}{3}\left[2\frac{\delta\chi}{\bar{\chi}}+\left(\frac{\delta\chi}{\bar{\chi}}\right)^{2}\right].$ (64)
B.3 Identifying $f_{NL}$
We decompose $\zeta$ into a Gaussian part $\zeta_{g}$ (linear term) and a non-Gaussian correction. From Eq. (64), the linear Gaussian part is:
$\zeta_{g}=\frac{2r_{dec}}{3}\frac{\delta\chi}{\bar{\chi}}.$ (65)
We can express the quadratic term $(\delta\chi/\bar{\chi})^{2}$ in terms of $\zeta_{g}$:
$\left(\frac{\delta\chi}{\bar{\chi}}\right)^{2}=\left(\frac{3}{2r_{dec}}\zeta_{g}\right)^{2}=\frac{9}{4r_{dec}^{2}}\zeta_{g}^{2}.$ (66)
Now, substituting this back into Eq. (64):
$\zeta=\zeta_{g}+\frac{r_{dec}}{3}\left(\frac{9}{4r_{dec}^{2}}\zeta_{g}^{2}\right)=\zeta_{g}+\frac{3}{4r_{dec}}\zeta_{g}^{2}.$ (67)
The standard definition of the local non-linearity parameter $f_{NL}$ is given by the expansion:
$\zeta=\zeta_{g}+\frac{3}{5}f_{NL}\zeta_{g}^{2}.$ (68)
Comparing the coefficients of the $\zeta_{g}^{2}$ term:
$\frac{3}{5}f_{NL}=\frac{3}{4r_{dec}}\quad\Longrightarrow\quad f_{NL}=\frac{5}{4r_{dec}}.$ (69)
Appendix C Tensor Perturbations with Non-minimal Coupling
In this appendix, we verify the stability of the tensor perturbations and estimate the corrections to the primordial gravitational waves induced by the non-minimal coupling term $\xi\chi^{2}R$.
C.1 Second-Order Action for Tensor Modes
We consider the tensor perturbations to the metric in the Jordan frame, defined as $g_{ij}=a^{2}(\tau)(\delta_{ij}+h_{ij})$, where $h_{ij}$ satisfies the transverse-traceless conditions ($\partial^{i}h_{ij}=0$, $h^{i}_{i}=0$). The second-order action for tensor modes arising from the gravitational sector of Eq. (1) is given by:
$S_{h}=\frac{1}{8}\int d\tau d^{3}x\,a^{2}\Omega(\chi)\left[(h_{ij}^{\prime})^{2}-(\nabla h_{ij})^{2}\right],$ (70)
where primes denote derivatives with respect to conformal time $\tau$. The non-minimal coupling modifies the effective Planck mass squared:
$M_{\rm eff}^{2}(\chi)\equiv\Omega(\chi)=M_{P}^{2}-\xi\chi^{2}.$ (71)
C.2 Evolution Equation
To diagonalize the kinetic term, we define the canonical variable $v_{k}(\tau)=z_{T}h_{k}(\tau)$ for each polarization mode, with the pump field defined as:
$z_{T}(\tau)=\frac{a(\tau)}{2}\sqrt{\Omega(\chi)}.$ (72)
The equation of motion (Mukhanov-Sasaki equation) for the mode function $v_{k}$ in Fourier space is:
$v_{k}^{\prime\prime}+\left(k^{2}-\frac{z_{T}^{\prime\prime}}{z_{T}}\right)v_{k}=0.$ (73)
During slow-roll inflation, the curvaton field $\chi$ is effectively frozen (or evolving very slowly compared to the expansion), so $\Omega(\chi)\approx\text{const}$. In the de Sitter limit where $a(\tau)\approx-1/(H\tau)$, the effective mass term simplifies to:
$\frac{z_{T}^{\prime\prime}}{z_{T}}\approx\frac{a^{\prime\prime}}{a}\approx\frac{2}{\tau^{2}}.$ (74)
This indicates that the evolution of tensor modes remains effectively the same as in General Relativity, up to a rescaling of the normalization factor $z_{T}$.
C.3 Primordial Power Spectrum
Solving the mode equation with Bunch-Davies vacuum initial conditions leads to the primordial tensor power spectrum on super-horizon scales:
$\mathcal{P}_{T}(k)=\frac{2H^{2}}{\pi^{2}M_{\rm eff}^{2}}=\frac{2H^{2}}{\pi^{2}(M_{P}^{2}-\xi\chi_{*}^{2})},$ (75)
where quantities are evaluated at the horizon exit. We can express this relative to the standard GR prediction $\mathcal{P}_{T,GR}\approx\frac{2H^{2}}{\pi^{2}M_{P}^{2}}$:
$\mathcal{P}_{T}(k)=\mathcal{P}_{T,GR}\left[1-\xi\left(\frac{\chi_{*}}{M_{P}}\right)^{2}\right]^{-1}\approx\mathcal{P}_{T,GR}\left[1+\xi\frac{\chi_{*}^{2}}{M_{P}^{2}}\right].$ (76)
C.4 Consistency Check: Sub-dominance Condition
The validity of our model relies on the assumption that the curvaton $\chi$ is a spectator field during inflation, meaning its energy density is negligible compared to the inflaton, and its field value is sub-Planckian:
$\chi_{*}\ll M_{P}\quad\Longrightarrow\quad\frac{\chi_{*}^{2}}{M_{P}^{2}}\ll 1.$ (77)
Assuming the field does not traverse trans-Planckian distances from inflation to today, the correction term $\xi(\chi/M_{P})^{2}$ remains negligible throughout cosmic history.
Appendix D Appendix: Derivation of the Effective Equation of State
We consider the Running Curvaton model extended with a non-minimal gravitational coupling in the Jordan Frame. The action is given by (1), where the non-minimal coupling function is defined as:
$\Omega(\chi)=M_{P}^{2}-\xi\chi^{2}.$ (78)
Then, we will varying with respect to this action for the gravitational sector, matter sector and dark energy sector, respectively. First, let us varying with the gravitational sector $\delta S_{g}=\frac{1}{2}\int d^{4}x\delta\left(\sqrt{-g}\Omega(\chi)R\right)$. Essentially, we need to deal with
$\delta(\sqrt{-g}\Omega R)=(\delta\sqrt{-g})\Omega R+\sqrt{-g}\Omega(\delta R)$
$=-\frac{1}{2}\sqrt{-g}g_{\mu\nu}\Omega R\delta g^{\mu\nu}+\sqrt{-g}\Omega(R_{\mu\nu}\delta g^{\mu\nu}+g_{\mu\nu}\Box\delta g^{\mu\nu}$ (79)
$-\nabla_{\mu}\nabla_{\nu}\delta g^{\mu\nu}).$
Since $\Omega(\chi)$ is not a constant, the total divergence term in $\delta R$ cannot be discarded. We perform integration by parts to shift the derivatives onto $\Omega$:
$\int d^{4}x\sqrt{-g}\Omega(g_{\mu\nu}\Box-\nabla_{\mu}\nabla_{\nu})\delta g^{\mu\nu}=$
$\int d^{4}x\sqrt{-g}(g_{\mu\nu}\Box\Omega-\nabla_{\mu}\nabla_{\nu}\Omega)\delta g^{\mu\nu}.$ (80)
Combining these terms, the variation of the gravitational sector is:
$\frac{\delta S_{g}}{\delta g^{\mu\nu}}=\sqrt{-g}\left[\frac{1}{2}\Omega(R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R)+\frac{1}{2}(g_{\mu\nu}\Box\Omega-\nabla_{\mu}\nabla_{\nu}\Omega)\right].$ (81)
Nextly, we will vary with respect to the curvaton part. This is the standard result for a canonical scalar field:
$\frac{\delta S_{\chi}}{\delta g^{\mu\nu}}=\frac{\delta}{\delta g^{\mu\nu}}\int d^{4}x\sqrt{-g}\left[\frac{1}{2}\Omega R-\frac{1}{2}g^{\mu\nu}\nabla_{\mu}\chi\nabla_{\nu}\chi-V\right]$ (82)
$=$ $\left(-\frac{1}{2}\sqrt{-g}g_{\mu\nu}\right)\left(-\frac{1}{2}\nabla_{\rho}\chi\nabla^{\rho}\chi-V\right)+\sqrt{-g}\left(-\frac{1}{2}\nabla_{\mu}\chi\nabla_{\nu}\chi\right)$
$=$ $\sqrt{-g}\left(-\frac{1}{2}\nabla_{\mu}\chi\nabla_{\nu}\chi+\frac{1}{4}g_{\mu\nu}(\nabla\chi)^{2}+\frac{1}{2}g_{\mu\nu}V(\chi)\right).$
Combined with the above results, we could derive the modified Einstein equation under the Jordan frame as follows,
$\Omega(\chi)G_{\mu\nu}=T_{\mu\nu}^{(m)}+T_{\mu\nu}^{(\chi)}+\nabla_{\mu}\nabla_{\nu}\Omega(\chi)-g_{\mu\nu}\Box\Omega(\chi),$ (83)
where $T_{\mu\nu}^{(\chi)}=\nabla_{\mu}\chi\nabla_{\nu}\chi-g_{\mu\nu}[\frac{1}{2}(\nabla\chi)^{2}+V(\chi)]$ is the energy-momentum tensor of the scalar field.
In this work, we will adopt the conformal flat FRW metric (3). In light of this background metric, the non-vanishing elements for the modified Einstein equations are summarized as
$G_{00}=3H^{2},G_{ij}=-a^{2}\eta_{ij}(-3H^{2}-2\dot{H}),$ (84)
Keeping in mind that $\chi$ only depends on the time. Thus, we have $\Box\Omega=g^{\mu\nu}\nabla_{\mu}\nabla_{\nu}\Omega=-\ddot{\Omega}-3H\dot{\Omega}$ and $\nabla_{0}\nabla_{0}\Omega=\ddot{\Omega}$, $\nabla_{i}\nabla_{j}\Omega=-\Gamma^{0}_{ij}\dot{\Omega}=-a^{2}H\delta_{ij}\dot{\Omega}$.
Taking these calculations into account, we first could derive the $00$ component of the modified Einstein equation (83) as follows,
$3(M_{P}^{2}-\xi\chi^{2})H^{2}=\rho_{m}+\frac{1}{2}\dot{\chi}^{2}+V(\chi)+6\xi H\chi\dot{\chi}.$ (85)
Similarly, we could derive the $ij$ component for Einstein equation as
$\Omega(\chi)\left(-2\dot{H}-3H^{2}\right)=p_{m}-\left(-\frac{1}{2}\dot{\chi}^{2}+V(\chi)\right)$-
$\left(2\xi\left(\dot{\chi}^{2}+\chi\ddot{\chi}\right)+4\xi H\chi\dot{\chi}\right)$ (86)
During these calculations, we have utilized
$\dot{\Omega}$ $=-2\xi\chi\dot{\chi},$ (87)
$\ddot{\Omega}$ $=-2\xi(\dot{\chi}^{2}+\chi\ddot{\chi}).$ (88)
We recast the modified Friedmann equations into the standard Einsteinian form to define effective dark energy quantities:
$3M_{P}^{2}H^{2}$ $=\rho_{m}+\rho_{DE}^{eff},$ (89)
$-2M_{P}^{2}\dot{H}$ $=(\rho_{m}+p_{m})+(\rho_{DE}^{eff}+p_{DE}^{eff}).$ (90)
According to these two definitions, we could easily obtain the effective energy density (27) and pressure (28) for DE induced by the curvaton.